# Interpret the Internal States of Recommendation Model with Sparse Autoencoders

### v2

@@ 5.3. Reconstruction and Interpretation Frontier with Scale Size $s$ and Sparsity Factor $k$ (Q3)
Figure 4. RecSAE achieves optimal reconstructions when the scale size $s$=16, corresponding to 1024 latents (16* 64 base model dimensions). This trend is consistent for both sparsity factor $k$=8 and $k$=16; in the figure, results are shown for $k$=8.
We conduct a series of experiments to examine the impact of hyper-parameters. Scale size $s$ determines latent space dimensions and sparsity factor $k$ controls the sparsity level of latent activations.
Scale Size $s$. We consider the reconstruction performance when choosing RecSAE scale sizes. We first fix the sparsity factor $k$=8 and vary $s$. As shown in Figure 4, reconstruction performance peaks when $s$=16, corresponding to 1024 latent dimensions. This suggests that 1024 latents are sufficient to encode the 64-dimensional dense internal representations of the investigated recommendation models. This pattern holds for both $k$=8 and $k$=16. Based on these observations, we set $s$=16 in experiments.
@@
Sparsity Factor $k$. $k$ controls the number of active latents in the latent space, impacting both the reconstruction quality and the interpretability of the RecSAE framework. Figure 5 illustrates the trade-off of the two factors. LightGCN is not included in the figure as its pattern is similar to BPRMF. It is important to note that RecSAE does not affect model performance, as the reconstructions do not need to replace the original model activations. Instead, it measures training performance of RecSAE to ensure that the latent space encodes the main information of the interpreted model. As shown in the figure, increasing $k$ typically improves reconstruction performance but results in fewer concepts with high interpretation confidence. As $k$ increases, each item/user is encoded by more latents, leading to a more redundant and less semantically distinct latent space. However, in two cases (BPRMF and TiMiRec on Amazon), the curve peaks at $k=8$, likely because smaller values of $k$ may lead to a relatively small number of activations, causing some latents to be insufficiently explained. In experiments, we set $k$=8 as a balance point between reconstruction quality and interpretability.
5.4. Training Effectiveness (Q4)
We further evaluate the effectiveness of RecSAE’s training by measuring reconstruction performance using MSE and assessing downstream recommendation performance on the test dataset. Table 6 presents the results on the Amazon dataset, with the other two datasets showing similar patterns. The notable reduction in MSE losses and significant improvements in recommendation performance from initialization to after training confirms the effectiveness of RecSAE’s training design. These strong reconstruction performances demonstrate that RecSAE latents effectively capture key information from the original model’s activations, while the reconstruction itself is not required during model interpretation.
Table 6. RecSAE Reconstruction Performance before → after training on Amazon, validating the training effectiveness.
Base Model MSE ↓ HR@10 ↑ NDCG@10 ↑
BPRMF 2.1235 → 0.1331 0.0077 → 0.0724 0.0033 → 0.0371
LightGCN 1.5552 → 0.2178 0.0029 → 0.0413 0.0013 → 0.0212
SASRec 2.0029 → 0.1949 0.0111 → 0.1534 0.0051 → 0.0830
TiMiRec 1.4121 → 0.2656 0.0042 → 0.1576 0.0019 → 0.0870
5.5. Potentials for Targeted Debiasing (Q5)
The RecSAE framework has the potential to make predictable changes to model behavior. Table 7 present case studies showing that decreasing and increasing latent activations leads to corresponding model output behavior. Furthermore, RecSAE may uncover concepts related to popularity bias or unfairness toward users and items. By revising these ”harmful” latents with tuned fairer reconstructions, recommenders can be effectively de-biased.
Table 7. Targeted Tuning. We scale latent value and examine whether the concept appears in top 10 recommended items.
Dataset Latent ID Concept Hit@10
-10× Original +10×
Amazon 288 Coconut 0.0189 0.6604 0.9528
MovieLens 463 Horror 0.0000 0.9213 1.0000
LastFM 67 Taylor Swift 0.0000 0.4079 1.0000
@@ 6. Discussion and Future work
Complexity. Assuming the probing site in the original model has a dimension of $|D|$ (typically less than hundreds in recommenders), the space complexity is $O(s|D|^{2})$, where $s$ is the scale size of latent space (in our experiments the optimal $s$ is 16). Training and inference have a time complexity of $O(s|D|^{2})$. As recommendation models typically do not scale with embedding dimensions (Sethi et al., 2022), and the complexity of RecSAE depends only on the scale size and embedding dimensions of the interpreted model, rather than the overall complexity of the base recommender, the RecSAE framework holds the potential for interpreting industry-level recommendationers.
Future Work. 1) Generalize to more types of recommendation models, such as content-based models and industry-level systems; 2) Investigate the impact of different probing positions across model layers to better understand information processing dynamics; 3) Conduct detailed evaluations of RecSAE’s utility for targeted tuning and debiasing to improve model fairness.
@@ 7. Conclusion
In this work, we propose the RecSAE framework with a probing paradigm for interpreting various types of recommendation models. It extracts interpretable latents from models’ internal activations and links these latents to semantic concepts, which are further validated to obtain confidence scores. These concepts are human-readable and easy to verify. The RecSAE framework represents the first step toward interpreting models without compromising their performances, while also enabling targeted tuning to models based on the interpretation results. Experiments conducted on three types of recommendation models across widely used datasets demonstrate the effectiveness and generalizability of the RecSAE framework.

### v3

@@ 7.1. Scale Size $s$
Figure 7. RecSAE achieves optimal reconstructions when the scale size ss=16, corresponding to 1024 latents (16* 64 base model dimensions). This trend is consistent for both sparsity factor kk=8 and kk=16; in the figure, results are shown for kk=8.
We first investigate the impact of varying the scale size $s$ on reconstruction performance. In this setting, we fix the sparsity factor at $k=8$ and vary $s$ to explore how the dimensionality of the latent space influences RecSAE’s ability to reconstruct the internal representations of recommendation models.
As shown in Figure 7, in most settings (3 out of 4), the reconstruction performance peaks at $s=16$, which corresponds to 1024 latent dimensions. Deviating from this configuration, either by reducing or increasing the number of latent dimensions, generally results in a degradation of reconstruction quality. This finding suggests that a latent space of moderate size is sufficient to capture the essential information encoded in the 64-dimensional dense representations output by the investigated models. Based on these results, we adopt $s=16$ as the default configuration in our experiments.
@@ 7.2. Sparsity Factor $k$
$k$ controls the number of active latents in the latent space, impacting both the reconstruction quality and the interpretability of the RecSAE framework. Figure 8 illustrates the trade-off of the two factors. LightGCN is not included in the figure as its pattern is similar to BPRMF. It is important to note that RecSAE does not affect model performance, as the reconstructions do not need to replace the original model activations. Instead, it measures training performance of RecSAE to ensure that the latent space encodes the main information of the interpreted model. As shown in the figure, increasing $k$ typically improves reconstruction performance but results in fewer concepts with high interpretation confidence. As $k$ increases, each item/user is encoded by more latents, leading to a more redundant and less semantically distinct latent space. However, in two cases (BPRMF and TiMiRec on Amazon), the curve peaks at $k=8$, likely because smaller values of $k$ may lead to a relatively small number of activations, causing some latents to be insufficiently explained. In experiments, we set kk=8 as a balance point between reconstruction quality and interpretability.
8. Targeted Tuning
In the preceding section, we demonstrated the efficacy of RecSAE in interpreting various types of recommendation models. Building upon these results, this section explores the potential of RecSAE in enabling targeted tuning of model behavior based on the interpretation outcomes. Specifically, the tuning is achieved by modifying the latent activations, which are then reconstructed and used to replace the original recommendation representations. As a result, the modified reconstructions either amplify or attenuate the targeted information.
We begin by presenting case studies where we vary the activation scale of latent variables. This variation leads to monotonic trends in the concept presence ratio among recommended items. Subsequently, we quantitatively report the impact of altering each latent variable and average concept presence ratios, thereby providing evaluations of the overall performance.
Formally, the concept presence ratio is calculated as follows, where Test represents the entire test set, $\mathbf{S}_{l}^{(t)}$ denotes the activation value of latent $l$ for input case $t$, $\mathbf{C}_{l}$ represents the interpreted concept associated with latent $l$, and $I_{pred}^{(t)}$ is the predicted item list (Top 10 items in later evaluations) for input case $t$. The indicator function $\mathbb{I}(\cdot)$ is used to evaluate whether a given condition holds true.
(5) $\begin{split}\text{Presence Ratio}(l)=\frac{\sum_{t}^{\text{Test}}\mathbb{I}\left(\mathbf{S}_{l}^{(t)}\neq 0\land\mathbf{C}_{l}\in i_{pred}^{(t)}\right)}{\sum_{t}^{\text{Test}}\mathbb{I}\left(\mathbf{S}_{j}^{(t)}\neq 0\right)}\end{split}$
8.1. Case Study
As shown in Table 6, RecSAE maps distinct item features to separate latents, demonstrating the model’s ability to isolate and represent individual features. Building on this, we identify these latents for targeted tuning. To explore this capability, we manipulate the activation of a single latent while keeping all others fixed, and observe the resulting changes in the model’s recommendation outputs.
Figure 9. By scaling the activation of each latent and measuring the resulting concept presence ratio in the recommended items, we demonstrate that this targeted modulation effectively steers the model’s outputs toward the corresponding direction.
In this process, the original activation values, denoted as scale 1, are systematically replaced with a set of scaled values: -10, -5 to 5 in steps of 1, and 10. For each of these scales, we obtain the corresponding recommendation outputs of the model, and compute the concept Presence Ratio, as defined in Equation 5. The results, presented in Figure 9, reveal a clear trend: increasing the activation values of latents leads to a higher ratio of items related to the target concept in the recommendations, while decreasing the activation values results in a reduction in the presence of such items. These monotonic trends illustrate RecSAE’s capacity for targeted tuning, offering a robust mechanism for adjusting model outputs based on specific latent activations.
In practice, this mechanism enables the amplification of particular concepts, such as child-friendly categories (e.g., “children’s,” “animation”), or the attenuation of undesirable content for children (e.g., “crime,” “thriller”) in the model’s outputs. This provides a powerful tool for targeted tuning model outputs to align with certain desired preferences and requirements.
8.2. Quantitative Report of Tuning Performance
The case study presented above demonstrates the monotonic relationship between latent activation scales and the corresponding concept presence in the output items. Building on these observations, we conduct further quantitative evaluations to assess the overall performance of tuning each latent. Specifically, we adjust the activation levels of each latent from a -10× scale to a +10× scale, obtaining the corresponding output recommendations. For each latent, we calculate the presence ratios of both the related and other concepts, reporting the averages across all latents. The presence ratio of the related concept reflects the efficacy of the tuning, and the presence ratio of the other latents serves as an indicator of the tuning precision, ensuring that unrelated concepts are minimally impacted.
The results are reported in Table 8. Across the three datasets, the monotonic trend of related concept persists: decreasing the activation values of the latents results in a lower concept presence ratio, while increasing the activation values leads to a higher ratio. The average variance in concept presence exceeds 10%. Although increasing the latent activations does not result in a 100% presence ratio, this outcome is understandable, given that there are eight latents activated in each cases and the modified latents may not dominate the overall behavior. If the original activation value of a latent is relatively low compared to others, scaling it may not always lead to the related concepts ranked in the top 10 item list.
Regarding the unrelated concepts, modifications across the three settings (decreased (-10×), original (+1×), and increased (+10×)) lead to minimal changes in outputs. s the rank of related items increases, some minor changes in the ratio of unrelated items are observed; however, these shifts remain minimal, not exceeding 0.5%. This is significantly lower than the changes observed in the related concepts. The relatively small variation in the ratio of unrelated items, in contrast to the more substantial shifts in related item presence, highlights the precision and effectiveness of RecSAE’s targeted tuning.
In conclusion, RecSAE demonstrates its potential as a powerful tool for targeted tuning of model behavior. By carefully adjusting the activation values of relevant and irrelevant latents, and applying these tuned reconstructions, recommendation models can be effectively customized for specific tasks and requirements.
Table 8. Targeted Tuning to SASRec trained on the three dataset. we modify the activation value of each latent by scaling it to -10× and +10×, and report the presence ratios for Related Concept and Other Concepts under the three settings: Decreased (-10×), Original (+1×), and Increased (+10×). The high variation in related concepts and the minimal change in other concepts demonstrate the efficacy of RecSAE’s targeted tuning capabilities.
Dataset Related Concept Other Concepts
Decreased Original Increased Decreased Original Increased
Amazon 9.3% 30.1% 43.2% 6.7% 7.2% 6.9%
MovieLens 3.8% 35.7% 47.5% 2.4% 2.9% 2.7%
LastFM 11.2% 23.3% 32.3% 4.3% 4.0% 4.0%
Finding 7: RecSAE shows the potential for targeted tuning of model behaviors.
@@ 9.1. Complexity
Assuming the probing site in the original model has a dimension of $|D|$ (typically less than hundreds in recommenders), the space complexity is $O(s|D|^{2})$, where $s$ is the scale size of latent space (in our experiments the optimal $s$ is 16). Training and inference have a time complexity of $O(s|D|^{2})$. As recommendation models typically do not scale with embedding dimensions (Sethi et al., 2022), and the complexity of RecSAE depends only on the scale size and embedding dimensions of the interpreted model, rather than the overall complexity of the base recommender, the RecSAE framework holds the potential for interpreting industry-level recommendation systems.
9.2. Future Work
This paper presents a structured design and evaluation of RecSAE’s capabilities. While RecSAE introduces a new paradigm for interpreting and tuning various recommendation models, several aspects remain open for further exploration. The key areas include:
1) Generalize to more types of recommendation models
RecSAE’s framework is highly flexible and scalable. The probing paradigm is adaptable, and the interpretation process is automated. These features make it suitable for integration with various types of recommendation models, including general, graph-based, context-based, sequential, content-based models and large-scale industry systems. Besides, industrial-scale recommendation systems often involve complex multi-stage pipelines (retrieval, ranking, and re-ranking). RecSAE can be integrated into such settings to evaluate how different stages capture varying levels of feature granularity—where the retrieval stage tends to model coarse-grained features, while finer-grained and more personalized features are progressively incorporated in the ranking and re-ranking stages. Future work could extend RecSAE to these additional models and stages, providing interpretations and comparable analysis to various type of systems.
2) Investigate the impact of different probing positions across model layers
As user embeddings are present in almost all recommendation models and serve as the core representation for personalized modeling, we probe this layer to evaluate the generalizability and effectiveness of RecSAE across diverse backbones. We acknowledge that model-aware probing at deeper layers reflecting the model’s internal reasoning process would provide complementary insights. Future research could explore the impact of probing at different positions within the model layers, based on the specific architecture of each model. This could provide deeper insights into how information is processed across layers and help researchers and system developers better understand the internal workflows of recommendation models. Additionally, investigations could also examine the concept formulation dynamics across model training process. This could reveal how the model progressively learns different levels of concepts as training evolves.
3) Establish more fine-grained and in-depth training and interpretation pipelines
The current two-step paradigm, comprising unsupervised training followed by post-hoc interpretation, offers certain advantages. It allows concepts to emerge objectively from the learned latents, without relying on a pre-defined concept set that could introduce bias during alignment training. Future directions could also explore a supervised method during training with alignment losses for better interpretation performance. For interpretation, the current pipeline consists of two components: concept construction and concept verification. It also employs two tools: a TF-IDF-based method for lightweight concept extraction and an LLM for deeper semantic understanding. While the current designs are carefully established, future work could explore more fine-grained pipelines, such as incorporating n-gram features into the TF-IDF stage, penalizing terms appearing in negative samples during construction, utilizing stronger LLMs beyond the current 8B open-source model, and developing more robust verification procedures. Moreover, future work could incorporate uncertainty-aware interpretation methods, such as selecting items based on a lower confidence bound to identify those most strongly associated with each latent’s semantics, rather than directly relying on the Top-N recommended items for the most activated cases. These improvements could further enhance the quality of the interpreted concepts derived from the learned sparse latents.
4) Conduct detailed evaluations of RecSAE’s utility for targeted tuning and de-biasing
In this paper, we demonstrate the potential of RecSAE for targeted tuning of model behavior. Specifically, Section 8 shows how RecSAE can influence the presence of interpreted concepts in model outputs by modifying latent activations. Further developments could focus on using RecSAE for targeted debiasing to enhance model performance and fairness. For example, RecSAE could help reduce the sensitive model outputs to specific users, such as child-unfriendly categories, and promote more diverse or niche content. Additionally, RecSAE could be used to mitigate popularity bias in recommendation models, improving fairness. These capabilities highlight the promising potential of RecSAE to address issues and improve performance in recommendation systems.
@@ 10. Conclusion
In this work, we propose the RecSAE framework with a probing paradigm for interpreting various types of recommendation models. It extracts interpretable latents from models’ internal activations and links these latents to semantic concepts, which are further validated to obtain confidence scores. These concepts are human-readable and easy to verify. The RecSAE framework represents a novel step toward interpreting models without compromising their performances, while also enabling targeted tuning to models based on the interpretation results. Experiments conducted on three types of recommendation models across widely used datasets demonstrate the effectiveness and generalizability of the RecSAE framework.
Acknowledge
This work is supported by the Natural Science Foundation of China (Grant No. U21B2026, 62372260), the National Key Research and Development Program of China under Grant 2024YFC3307403, and Research Project of Quan Cheng Laboratory, China (Grant No. QCL20250105).